### v1

# Public Projects with Preferences and Predictions

Abstract
In the public projects problem, a group of decisionmakers aggregate their preferences to choose one alternative. Recent work on public projects has proposed the Quadratic Transfers Mechanism (QTM) and shown asymptotic welfare guarantees in some cases. We begin by giving new non-asymptotic Price of Anarchy guarantees for the QTM.
We then incorporate an alternative philosophy toward group decisionmaking, aggregation of information about which is the best alternative. We propose a public projects mechanism based on the QTM that aggregates both preferences and predictions, modeled as forecasts of the projects’ welfare impacts. When the predictions come from a prediction market or wagering mechanism, we show the entire mechanism is robust to manipulation and give Price of Anarchy guarantees, though under strong assumptions on the mechanism’s knowledge. Our results focus primarily on the case of deciding between two alternatives, showing the Price of Anarchy tends to $1$ as natural measures of the “size” of the population grow large. In most cases, the mechanisms achieve a balanced budget as well.
1 Introduction
In the public projects problem, a group of decisionmakers is faced with selecting one of multiple alternative projects to undertake. Classically, the objective is to aggregate the preferences of the decisionmakers in order to make a choice which most benefits the group as a whole. This paradigm includes social choice theory and mechanism design.
An alternative philosophy is to focus on aggregating information about which alternative is the best. For example, a committee could make a decision either by voting (preferences), or by discussing until rough consensus is reached (information). Decision markets [Othman and Sandholm, 2010] have been proposed as a group decisionmaking mechanism in the information paradigm. There, forecasts about the impact of each alternative are produced by running prediction markets – financial markets whose prices reflect predictions – and then a decision is automatically chosen based on the predicted-best alternative. We likewise focus on predictions as a source of credible, (partly) verifiable information.
A public-projects problem can generally benefit from aggregation of both preferences and information. Can we design a mechanism for groups to make decisions based on both? There are only a few formal proposals in this spirit (see Section 1.2). When agents can serve as both information aggregators and preference-driven decisionmakers, they may have incentives to manipulate in both roles.
In this paper, we will propose a mechanism that aggregates both preferences and predictions to make a group decision. For quantitative results, we consider monetary mechanisms and assume quasilinear utility, seeking social welfare guarantees. Each participant $i$ has a value $v_{k}^{i}\geq 0$ for each project $k$, with $V_{k}=\sum_{i}v_{k}^{i}$. Meanwhile, project $k$ has some unknown expected external welfare impact $B_{k}$, a function of participants’ private information. We seek to elicit the preferences $\{v_{k}^{i}\}$ and information $\{B_{k}\}$ and maximize social welfare, $\arg\max_{k}(V_{k}+B_{k})$. The Synthetic Players Quadratic Transfers Mechanism with Predictions (SQUAP) begins with an information aggregation phase, such as a prediction market or a wagering mechanism, to elicit $\{B_{k}\}$. Then, it uses a “voting” stage, based on the recently-introduced Quadratic Transfers Mechanism discussed below, to elicit preferences and select an alternative. The mechanism is practical and budget-balanced. Unfortunately, it is naturally also very difficult to analyze strategically. However, by adding some assumptions on the designer’s knowledge, we can obtain a less-practical variant we call Impractical SQUAP. Then we can prove the following bound on the Price of Anarchy, or worst-case equilibrium ratio of the mechanism’s welfare to optimal-achievable welfare.
Theorem (Main result, informal).
In Impractical SQUAP, using either a prediction market or wagering mechanism, for two alternatives,
Price of $\text{Price of Anarchy}\geq 1-O(T^{-2/5}),$
where the spread $T=\frac{\text{total welfare}}{\text{maximum individual value}}$ measures the size of the game.
More formally, $T=\frac{\arg\max_{k}(V_{k}+B_{k})}{\max_{i,k}v_{k}^{i}}$, the ratio of first-best welfare to the largest individual value. As the total welfare grows large relative to the maximium influence of any one person’s preference, the Price of Anarchy tends to one. This result proves that it is possible to align incentives when strategic agents are both forecasters and decisionmakers. We view the impractical mechanism’s welfare guarantee as strong evidence for the welfare of our main proposed mechanism.
1.1 Outline and Results
The QTM.
Recent work [Eguia et al., 2019, 2023] has introduced and studied the Quadratic Transfers Mechanism (QTM) for public projects. Inspired by quadratic voting [Lalley and Weyl, 2018, Goeree and Zhang, 2017], the mechanism collects bids (“votes”) on each alternative, with cost rising quadratically, and picks an alternative according to a probabilistic “softmax” of the vote totals. By redistributing payments, the mechanism is budget-balanced.
To analyze our combined aggregation-decision mechanisms, we first analyze the QTM and present Price of Anarchy guarantees (Section 2). Our approach and results build on the work of Eguia et al. [2019], which gives asymptotic welfare bounds in a model with agents drawn i.i.d. from bounded value distributions, where realized values are common knowledge. We consider any worst-case set of agents with no distributional assumptions. We analyze pure-strategy Nash equilibrium, a natural solution concept in this setting for several reasons:
  • Pure strategies are natural because the strategy space in the QTM is a convex set, $\operatorname{\mathbb{R}}^{m}$, the number of “votes” to cast on each alternative. Further, we show that for a reasonable setting of the QTM parameter, utilities are strictly concave, implying that if others play pure strategies, it is strictly suboptimal to randomize.
  • (Nash vs Bayes-Nash, I) In the QTM, computing a best-response only requires knowledge of opponents’ aggregate vote totals. The situation is reminiscent of mean field games [Caines, 2021]. We view this amount of knowledge as equally reasonable to a Bayes-Nash equilibrium assumption that participants know e.g. the distribution of the vote totals.
  • (Nash vs Bayes-Nash, II) A designer can facilitate the spread of information required to achieve Nash equilibrium, e.g. with straw poll, or a dynamic voting process that publicizes information about the vote totals during the mechanism. On the other hand, if a Bayes-Nash analysis were to rely on agents’ ignorance of the others’ realized types, it could be fragile.
Types of information.
Broadly, there are two types of information relevant to the decision: internal-relevant and external-relevant. Internal-relevant information affects decisionmakers’ own preferences. External-relevant information affects estimated “external welfare” of the decision, i.e. impacts on others. For example, consider shareholders of a company deciding between two product lines. The expected revenue of each product would be internal-relevant, as each shareholder’s utility and preference is related to the revenue. The climate impact (e.g. expected $\text{CO}_{2}$ emissions) would be external-relevant, as it does not impact each shareholder’s utility directly (for the sake of argument). We focus on modeling external-relevant information only, i.e. the external welfare impacts $\{B_{k}\}$. Our mechanism could easily be modified to elicit internal-relevant information in the first stage (e.g. with prediction markets). However, for the sake of analysis, we assume participant preferences $\{v_{k}^{i}\}$ are fixed independent of $\{B_{k}\}$.
External-relevant information is important in many public-projects or group decisionmaking contexts. If a governing board, shareholders, or co-op members are making a decision about their organization, their charter may require consideration of the impact on the organization as a whole, its employees, or the population the organization is intended to serve. This could hold particularly for a non-profit, co-op, or a Decentralized Autonomous Organization (DAO) in a blockchain-governance context [Spelliscy et al., 2024]. Such decisions often involve a conflict between what is best for the decisionmakers versus what is best for the organization’s mission. Our mechanism could help by aligning incentives and optimizing for a combination.
To move toward incorporating information, we consider in Section 3 an extension of the QTM to a case where the external welfare impact of each project is fixed and known. Our Synthetic Players QTM first collects agent votes, then “participates” with synthetic players whose values capture the external welfare impacts. Unfortunately, analysis is challenging. We consider an “impractical variant” that needs knowledge about the setting and commits to the synthetic votes first, then collects votes from the agents. In this case, we are again able to prove a strong Price of Anarchy bound for the two-alternative case.
Combined preference-prediction mechanisms.
In Section 4 we define a two stage mechanism, Synthetic QUAdratic Transfers with Predictions (SQUAP). In the first stage, agents participate in an information-aggregation mechanism, such as prediction market or wagering mechanism to estimate the external welfare impacts $\{B_{k}\}$ of the projects. Then, the second stage runs the Synthetic Players QTM from Section 3 with input $\{B_{k}\}$ to produce the final decision.
To prove a Price of Anarchy bound (for the impractical variant), we take the following steps. First, we consider two commonly-studied mechanisms for eliciting predictions: prediction markets and wagering mechanisms. Second, we show that, while agents may have an incentive to misreport predictions given the decision is based on their information, deviations will be bounded and tend to zero as the size of the information-aggregation mechanism grows. Finally, we use an importance-weighting technique from the decision markets literature [Chen et al., 2011] so that forecasters’ expected payoffs are independent of the decision of the mechanism. Under some mild assumptions, we provide a Price of Anarchy guarantee, our main result, based both on the spread of the values $T$ and the size of the information-aggregation mechanism.
1.2 Related Work
Public projects.
There is naturally a significant amount of work on public projects from a mechanism-design perspective. We focus on the setting of this paper, which is social welfare maximization with quasilinear-utility agents. A standard solution is to use the Vickrey-Clarke-Groves mechanism (VCG) [Vickrey, 1961, Clarke, 1971, Groves, 1973] which has an equilibrium that is truthful and maximizes social welfare. However, VCG has some undesirable properties, including lack of a Price of Anarchy bound (there exist arbitrarily bad equilibria) and uncertain revenue (it may be very large or zero). Literature addressing the problem with more nuance than our basic model considers the excludable and non-excludable cases, e.g. Ohseto [2000]. The smoothness framework has been used to show a constant-factor Price of Anarchy bound for a mechanism in which participants place bids on all alternatives, the one with the highest total is selected, and everyone pays their bid on the winner [Roughgarden et al., 2017].
QTM.
Recent research has considered transfer-based mechanisms which use quadratic payments  [Weyl, 2013, Lalley and Weyl, 2018, Goeree and Zhang, 2017]. Lalley and Weyl [2018, 2019] propose a binary-outcome quadratic voting mechanisms and show it is asymptotically efficient. Weyl [2017] then studies this mechanism’s sensitivity to collusion, fraud, and voter mistakes when the number of agents is large. A number of other works study the behavior or empirical performance of quadratic voting [Chandar and Weyl, 2019, Casella and Sanchez, 2019, Quarfoot et al., 2017, Goeree and Zhang, 2017].
Building on Lalley and Weyl [2019], the QTM in particular is proposed by Eguia et al. [2019] for $m\geq 2$ alternatives. The authors consider a setting where participants’ values for the $m$ alternatives are drawn i.i.d. from a distribution over $[0,1]^{m}$, and realized values are common knowledge. They prove that for an undetermined choice of the mechanism’s scale parameter $c$, equilibria exist, and show that the probability of selecting the highest-welfare alternative approaches 1 as the number of agents diverges. Eguia et al. [2023] then study the same mechanism in a similar setting but under Bayes-Nash mixed-strategy equilibrium, with similar results. The part of our work on the QTM builds heavily on the approaches in Eguia et al. [2019, 2023], but our results address the nonasymptotic regime without distributional assumptions. This includes concrete bounds with small constants.
Aggregating information for decisionmaking.
Many works study mechanisms for eliciting information from experts and using it to make a decision. The most relevant mechanisms for this paper are decision markets [Hanson, 1999, Othman and Sandholm, 2010], which simultaneously operate a prediction market for each alternative $k$. (A prediction market is a financial market designed to aggregate beliefs into a consensus forecast, reflected in the prices of the financial products being traded [Hanson, 2003].) The decision market mechanism selects the alternative that is best according to the aggregated predictions (or uses a similar rule). Even without preferences over the alternatives, decision markets can have complex incentive misalignment problems, although these can be fixed by the importance-weighting technique of Chen et al. [2011].
Consulting with experts.
There are many works in which a single decisionmaker elicits information from one or more experts prior to making a decision, and the experts wish to influence the decision. This includes the “cheap talk” model of Crawford and Sobel [1982] as well as Bayesian Persuasion [Kamenica and Gentzkow, 2011]. Oesterheld and Conitzer [2020] considers “decision scoring rules” that incentivize truthful predictions and recommendations by an expert.
For aggregating information from a group, e.g. Gerardi et al. [2009], consider a Bayesian setting in which a decisionmaker elicits information from a number of experts and aggregates it to make a decision. To incentivize truthfulness, the decisionmaker with small probability audits an expert’s report against the others (similar to peer prediction [Miller et al., 2005]) and picks an alternative desired by the audited agent.
Such settings typically do not face the concern in this paper, where the decisionmakers are a large public group and may overlap with the forecasters.
Decisions from preferences and predictions.
A few works are much closer to the spirit of our motivation, if technically quite different. They generally analyze voter preferences as a function of the information to be aggregated. In contrast, we model preferences as fixed with information as an orthogonal axis. In the “Wisdom-of-the-Crowd Voting Mechanism” of Schoenebeck and Tao [2021], each participant is both a voter and a holder of private information. The mechanism elicits both information and a preference in a single shot, then aggregates both to make a binary decision. The authors show that, with high probability, the mechanism selects despite strategic behavior the “majority wish” alternative that more voters would prefer if fully informed.
Amanatidis et al. [2022] consider a similar problem motivated by blockchain applications. Voters have private information and participate in an approval vote among $m\geq 2$ alternatives. As with Schoenebeck and Tao [2021], information aggregation takes place within the mechanism rather than through an explicit phase prior to voting. The authors prove that the mechanism can achieve a Price of Anarchy of $\frac{1}{2}$.
In Jackson and Tan [2013], a majority-vote between two alternatives takes place after a “deliberation” stage. The experts who take part in deliberation cannot also be voters (unlike in our model), and cannot misreport arbitrarily, but only choose to either reveal or hide their information in order to influence the voters. Other works in this spirit are Alonso and Câmara [2016], Schnakenberg [2015].
2 Quadratic Transfers Mechanism
In this section, we give results, including a Price of Anarchy bound, for the Quadratic Transfers Mechanism (QTM) for the public projects problem. These results complement existing analyses [Eguia et al., 2019, 2023], and will also serve as a foundation for our results in later sections combining predictions and preferences.
2.1 Model
Public projects.
We consider a classic public projects setting in which a group of $n$ agents needs to select one of $m\geq 2$ alternatives. We will use $i,j$ to denote generic agents and $k,l$ to denote generic alternatives. Let $v_{k}^{i}\in\operatorname{\mathbb{R}}_{\geq 0}$ represent agent $i$’s value for alternative $k$. We assume that all values are nonnegative. Denote $\mathbf{v}^{i}=(v_{1}^{i},v_{2}^{i},\ldots,v_{m}^{i})$ as the vector of agent $i$’s values across all $m$ alternatives. The aggregate value for alternative $k$ is $V_{k}:=\sum_{i=1}^{n}v_{k}^{i}$. Without loss of generality, we number the alternatives such that $V_{1}\geq V_{2}\geq\ldots\geq V_{m}$.
Mechanisms and welfare.
A mechanism defines a game played by the set of agents. Given the actions of the agents, the output of the mechanism is a probability distribution $\mathbf{p}\in\Delta_{m}$, the set of distributions over the alternatives, along with a net payment from each agent. We assume quasilinear utility: If the mechanism’s output is $\mathbf{p}$ and if agent $i$ makes a net payment $\pi_{i}$, then $i$’s expected utility is given by $\sum_{k}p_{k}v_{k}^{i}-\pi_{i}$. The social welfare of the mechanism is $\sum_{k}p_{k}V_{k}$, the expected aggregate value produced. The optimal or “first-best” social welfare is $V_{1}$, the aggregate value of the best alternative.
In general, the Price of Anarchy of a mechanism is the worst case ratio between the social welfare of the mechanism in equilibrium and the optimal social welfare. Formally, given a mechanism and a value profile $\mathbf{v}=(\mathbf{v}^{1},\dots,\mathbf{v}^{n})$, let $\text{pNE}(\mathbf{v})$ denote the set of distributions $\mathbf{p}\in\Delta_{m}$ induced by the mechanism in any pure-strategy Nash equilibrium. Then the pure-strategy Price of Anarchy of the game with respect to $\mathbf{v}$ is
$\text{pPoA}{(\mathbf{v})}=\frac{\min_{\mathbf{p}\in\text{pNE}(\mathbf{v})}\sum_{k}p_{k}V_{k}}{V_{1}}.$ (1)
We define the Price of Anarchy relative to a value profile $\mathbf{v}$ because we will give bounds depending on properties of $\mathbf{v}$; for instance, a Price of Anarchy bound that improves if all agents are “small”. However, for convenience, we may refer to the pPoA without explicitly including $\mathbf{v}$.
We say a mechanism achieves budget balance if the net total payment of all agents is zero.
The QTM.
In the Quadratic Transfers Mechanism (QTM) [Eguia et al., 2019, 2023], each agent $i$ simultaneously and privately submits a vector $\mathbf{a}^{i}\in\operatorname{\mathbb{R}}^{m}$, where the $k$th entry $a_{k}^{i}$ represents a number of “votes” for alternative $k$. Votes are allowed to be negative, representing voting against that alternative. We let $A_{k}:=\sum_{i}a_{k}^{i}$ denote the aggregate votes for alternative $k$.
Each agent is charged $c\sum_{k}(a_{k}^{i})^{2}$, i.e. the sum of squares of their votes multiplied by a parameter $c>0$ chosen in advance by the mechanism designer. Then, the mechanism chooses one alternative according to the “softmax” distribution $\mathbf{p}\in\Delta_{m}$ defined by
$p_{k}=\frac{e^{A_{k}}}{\sum_{l}e^{A_{l}}}.$
The QTM redistributes each agent’s payment equally among the other voters, so that $i$ receives
$\frac{c}{n-1}\sum_{j\neq i}\sum_{k}(a_{k}^{j})^{2}.$ (2)
Thus, the QTM is budget-balanced. [^1] This redistribution does not affect the incentives of each voter, so it can be modified without changing the strategic properties of the mechanism. We will utilize this in some of proposed mechanisms later in the paper. Agent $i$’s utility function $u^{i}$, taken in expectation over the mechanism’s randomness, as a function of the votes $\mathbf{a}\coloneqq(\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{n})$, is
$u^{i}(\mathbf{a})=\sum_{k=1}^{m}p_{k}v_{k}^{i}-c\sum_{k=1}^{m}(a_{k}^{i})^{2}+\frac{c}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}.$ (3)
We note that agent $i$’s actions only affect the first two terms, so for strategic analysis, the third (redistribution) term is irrelevant.
2.2 Equilibrium Characterization
In this subsection, we characterize conditions under which a pure-strategy equilibrium exists in the QTM. Our results build on some techniques of  Eguia et al. [2019], but also introduce new techniques and new structural features of the mechanism, such as the fact that votes across alternatives always sum to zero in equilibrium.
To show existence, we take the following steps. First (Lemma 7), we show that for each player, it is strictly dominated to play any $\mathbf{a}^{i}$ lying outside a certain compact set. It follows that all equilibria are “as if” the game were restricted to a compact strategy space (Lemma 8). From there, we can show that if the parameter $c$ is large enough relative to the maximum agent preference, then utilities are concave, and apply an existence theorem (Lemma 1). A similar argument appears in Eguia et al. [2019], but in our case we give a constructive bound on $c$ that appears quite tight. We also note a connection in the analysis to exponential family distributions. Finally, we derive the first-order conditions for pure-strategy Nash equilibrium, and show as a byproduct that each agent’s votes sum to zero.
All proofs appear in Appendix A.1.
Lemma 1.
In the QTM on agent value profile $\mathbf{v}$, if the mechanism chooses $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then every agent’s utility is strictly concave as a function of their vote $\vec{a}^{i}$, for any fixed strategy of their opponents.
Proposition 1.
In the QTM on agent value profile $\mathbf{v}$, if the mechanism chooses $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then a pure-strategy Nash equilibrium exists.
There may be many scenarios where a pure strategy equilibrium exists even for much lower values of $c$. For example, we conjecture this to be the case when agent values are “well-aligned” with each other, without much disagreement.
We now state the important first-order conditions that characterize equilibrium in terms of the mechanism’s expected individual and aggregate welfare, $\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i}=\sum_{k}p_{k}v_{k}^{i}$ and $\operatorname{\mathbb{E}}_{k\sim p}V_{k}=\sum_{k}p_{k}V_{k}$ respectively. This result also appears throughout Eguia et al. [2019], although we add some observations that are useful to us later in constructing equilibria as part of our mechanisms.
Lemma 2.
In any pure-strategy Nash equilibrium of the QTM with values $\mathbf{v}$, the votes $\mathbf{a}$ satisfy
$a_{k}^{i}$ $=\frac{p_{k}}{2c}\left(v_{k}^{i}-\sum_{\ell}p_{\ell}v_{\ell}^{i}\right)$ for all $i$, (4)
$A_{k}$ $=\frac{p_{k}}{2c}\left(V_{k}-\sum_{\ell}p_{\ell}V_{\ell}\right).$ (5)
Furthermore, if the mechanism sets $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then $\mathbf{a}$ is a pure-strategy Nash equilibrium if and only if (4) and (5) are satisfied. Furthermore, for any vote totals $\mathbf{V}$, if $\mathbf{A}$ is any aggregate vote vector that satisfies (5), then for any $\mathbf{v}$ summing to $\mathbf{V}$ there exists a corresponding $\mathbf{a}$ summing to $\mathbf{A}$ satisfying (4).
An equilibrium is unique if Equation 5 has a unique solution $\vec{A}$. We do not know conditions under which there is a unique pure-strategy equilibrium, but conjecture that it is often unique. Summing over Equations 4 and 5, we obtain:
Corollary 1.
The sum of votes for each agent across alternatives is 00 in equilibrium. The sum of aggregate votes across alternatives is 00 in equilibrium: $\sum_{k=1}^{m}A_{k}=0.$
2.3 Price of Anarchy: Two Alternatives
In this section, we present our results for the QTM on $m=2$ alternatives. Conceptually similar results have already been shown by Eguia et al. [2019, 2023] (see Section 1.2), but they focus more on the challenges of the $m\geq 3$ setting and have different distributional and model assumptions than we do here. Recall that $V_{1}$ is the aggregate value of the better alternative, i.e. optimal welfare, and $V_{2}$ is the aggregate value of the other. Our main result, Theorem 1, is that the QTM’s pure-strategy Price of Anarchy (pPoA) is always at least $\tfrac{1}{2}$ and approaches $1$ rapidly as the spread of the game — the ratio of $V_{1}$ to the largest individual value — grows. We also show in Corollary 2 a faster rate of convergence to $1$ when the gap — the difference between $V_{1}$ and $V_{2}$ — grows. Finally, we will characterize the total payments of the QTM in terms of the amount of “disagreement” of the participants.
Theorem 1.
Define the spread of the game to be $T\coloneqq\frac{V_{1}}{max_{i,k}v_{k}^{i}}$. In the two-alternative case, there exists a choice of $c$ such that the QTM achieves a Price of Anarchy
[…]
The result is proven in Appendix A.2. The proof relies on a careful analysis of first-order conditions. In particular, if the spread of the game is large, then there are two cases: $V_{2}$ is also large, in which case the welfare of any mechanism is large; or else the gap $V_{1}-V_{2}$ is large, in which case we can prove a pPoA bound. It also yields:
Corollary 2.
Define the gap of the game to be $G\coloneqq\frac{V_{1}-V_{2}}{max_{i,k}v_{k}^{i}}$. In the two-alternative case, there exists a choice of $c$ such that the QTM achieves a Price of Anarchy
[…]
Total payments.
Define the revenue of the QTM to be the total payments of the agents before redistribution, $c\sum_{i}\sum_{k}(a_{k}^{i})^{2}$. It is interesting, and will be useful later, to understand the relationship of revenue to the type profile $\mathbf{v}$. As with the pPoA analysis, we can analyze revenue using the first-order conditions.
Proposition 2.
In the QTM with two alternatives and type profile $\mathbf{v}$, with $V_{1}>V_{2}$, define the disagreement to be $D:=\frac{\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}}{(V_{1}-V_{2})^{2}}$. Then in any pure-strategy equilibrium, [^2] Here Θ()normal-$\Theta(\cdot)$ is with respect to $V_{1}-V_{2}$ growing, or at least bounded below. The proof includes fully explicit nonasymptotic bounds.
[…]
In other words, the revenue is tightly controlled by the disagreement $D$ of the strategy profile. For intuition, suppose the strengths of the agent preferences are identical with $|v_{1}^{i}-v_{2}^{i}|=1$. Then the numerator of $D$ is $n$. Let $n_{1}$ be the number of agents with $v_{1}^{i}-v_{2}^{i}=1$ and let $n_{2}$ be the number with $v_{1}^{i}-v_{2}^{i}=-1$. Observe that $V_{1}-V_{2}=n_{1}-n_{2}$, the number of agents who prefer alternative $1$ minus the number who prefer $2$.
  • If all agents agree, i.e. $n_{1}=n$, then $V_{1}-V_{2}=n$ and the denominator is $n^{2}$; revenue converges to zero quickly.
  • If $n_{1}-n_{2}=\Theta(\sqrt{n})$, then $D=\Theta(1)$, and revenue grows polylogarithmically in $n$.
  • For $n_{1}-n_{2}\ll\sqrt{n}$, $D=\omega(1)$ and revenue grows more rapidly.
The disagreement $D$ has a natural statistical interpretation. If we pick an agent $i$ uniformly at random and let $X=v_{1}^{i}-v_{2}^{i}$, then $D=\frac{1}{n}\left(\frac{\sigma^{2}+\mu^{2}}{\mu^{2}}\right)$, where $\mu$ and $\sigma$ are the mean and standard deviation of $X$.
2.4 $m>2$ Alternatives
We briefly touch on the case where there are more than two alternatives. Eguia et al. [2019, 2023] give asymptotic results showing that in their settings, with a fixed distribution of bounded agents, as $n\to\infty$ the welfare of the QTM tends to optimality. Here we can give a weaker nonasymptotic bound. The proof is not immediate and involves dividing the alternatives into a group with “high” welfare (higher than the mechanism’s expectation) and the remainder, an idea from Eguia et al. [2023] used in a different way.
Proposition 3.
[…]
It is not clear if this bound can be improved non-asymptotically. The analysis of the many-alternatives case is very complex (see Eguia et al. [2019, 2023]). As far as we currently know, it may be possible to have instances of a large “spread” $\frac{V_{1}}{\max_{i,k}v_{k}^{i}}$, perhaps exponentially large in $m$, with pure-strategy Price of Anarchy bounded away from one, or perhaps as small as $O(1/m)$.
3 Incorporating External Welfare
In this section, we propose mechanisms for public projects with external welfare, meaning estimates or projections of the impact of the project additional to the values of the agents participating in the mechanism. First, we will suppose that the external welfare impacts are fixed and known to the mechanism. In Section 4, we will use this section’s mechanism for our main problem, the case where external welfare impacts are forecasts from (manipulable) mechanisms.
3.1 Model and Mechanism
We define the problem of Public Projects with External Welfare, extending the public projects setting defined in Section 2. In addition to the agent values $\mathbf{v}=(\mathbf{v}^{1},\dots,\mathbf{v}^{n})$, for each $k\in\mathbb{Y}$ we now let $B_{k}\in\operatorname{\mathbb{R}}_{\geq 0}$ represent the impact to external welfare conditional on the $k$th alternative being chosen.
Given $\mathbf{B}\coloneqq(B_{1},\dots,B_{k})$, our mechanism will run a modified QTM to choose an alternative. Our goal is to choose an alternative which maximizes total welfare, represented by the sum $V_{k}+B_{k}$. Thus let $W_{k}\coloneqq V_{k}+B_{k}$, and assume without loss of generality that alternatives are numbered according to $W_{1}\geq W_{2}\geq\ldots\geq W_{m}$. Therefore, the optimal achievable welfare is $W_{1}$ and the welfare of a mechanism with output $\mathbf{p}$ is $\sum_{k}p_{k}W_{k}=\sum_{k}p_{k}(V_{k}+B_{k})$. We let $\text{pPoA}(\mathbf{v},\mathbf{B})$ be the pure-strategy Price of Anarchy of our mechanism on profile $(\mathbf{v},\mathbf{B})$.
Mechanism.
We propose the Synthetic Players Quadratic Transfer Mechanism for public projects with external welfare. The mechanism is inspired by the approach of creating a set of synthetic players whose preferences reflect the external welfare, then playing on their behalf in the QTM alongside the real participants. However, the mechanism itself is more concrete and based on the analysis of equilibrium conditions for the QTM.
We define two versions of the mechanism. The first is a practical variant that, unfortunately, we are unable to analyze. The second is a less-practical variant that assumes the mechanism has significant knowledge about the agent values. We give Price of Anarchy results for the second variant and conjecture that similar bounds hold for the first.
The mechanisms use the following equations adapted from the first-order conditions (Lemma 2), where $p_{k}=e^{A_{k}}/\sum_{\ell}e^{A_{\ell}}$.
$A_{k}$ $=p_{k}\left(W_{k}-\sum_{\ell}p_{\ell}W_{\ell}\right)$ $(\forall k),$ (6)
$a_{k}^{\text{mech}}$ $=p_{k}\left(B_{k}-\operatorname{\mathbb{E}}_{k\sim p}B_{k}\right)$ $(\forall k).$ (7)
Equation 6 gives the first-order conditions for the aggregate vote totals $A_{1},\dots,A_{m}$ in pure-strategy equilibrium when the total welfare of each alternative is given by $W_{1},\dots,W_{m}$. Equation 7 gives the best responses $a_{1}^{\text{mech}},\dots,a_{m}^{\text{mech}}$ to a particular aggregate total $A_{1},\dots,A_{k}$ by an agent with value vector $\mathbf{B}$. The equations generally form a fixed-point problem.
Synthetic Players QTM.
The Synthetic Players QTM proceeds as follows.
  1. 1.
    Set $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$.
  2. 2.
    Collect votes $\mathbf{a}^{1},\dots,\mathbf{a}^{n}$ from the agents.
  3. 3.
    Collect payments based on votes and redistribute as in the QTM, i.e. Equation 2.
  4. 4.
    Define $A_{k}=a_{k}^{\text{mech}}+\sum_{i}a_{k}^{i}$ and solve Equations 6 and 7 for $a_{1}^{\text{mech}},\dots,a_{m}^{\text{mech}}$ and $A_{1},\dots,A_{m}$.
  5. 5.
    Draw an alternative from $\mathbf{p}$ defined by $p_{k}=e^{A_{k}}/\sum_{\ell}e^{A_{\ell}}$.
The idea of the mechanism is to simulate a set of synthetic players whose total values are $\mathbf{B}$ and total votes are $\mathbf{a}^{\text{mech}}$. (The payments of the synthetic players are irrelevant and can be set to zero, as shown in our analysis of the second variant below.) In an equilibrium of the original QTM, given the votes of the real agents, the synthetic agents’ best responses and the QTM output are given by Equations 6 and 7.
Difficulty of analysis.
The catch is that the Practical Synthetic Players QTM gives the (real) agents commitment power. In an equilibrium of the original QTM, if a real agent deviates to another strategy, the other agents would continue to play the equilibrium. But in this mechanism, the synthetic agents calculate their strategies in response to the strategies of the real agents. This opens the possibility for manipulation. To see this, note that in the two-alternative case, a succinct description of the mechanism is to collect $\mathbf{a}$, then solve the fixed-point problem
$p_{1}=\frac{e^{\left(\sum_{i}a_{1}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{1}-B_{2})}}{e^{\left(\sum_{i}a_{1}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{1}-B_{2})}+e^{\left(\sum_{i}a_{2}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{2}-B_{1})}},$ (8)
with $p_{2}=1-p_{1}$. We conjecture this mechanism has roughly the same Price of Anarchy guarantees as the impractical variant below. However, the change has an undeniable strategic impact. A tempting modification would be to replace each $p_{1}p_{2}$ in (8) with an estimated constant, such as $\tfrac{1}{4}$, removing the ability of agents to influence the impact of the external welfare terms and replacing the fixed-point problem with a simple equation. However, for that modification, there exist examples with low-welfare equilibria in which the influence of $\mathbf{B}$ dominates the decision, although $\mathbf{V}$ has more importance.
Synthetic Players QTM – Impractical Variant.
We now propose a variant that is less practical, but has strong theoretical guarantees. Specifically, we assume the mechanism has knowledge of $\mathbf{V}=(V_{1},\dots,V_{m})$, the total values of the agents. (The mechanism does not require knowledge of individual values beyond an upper bound on the maximum, as in the QTM.) With this, we flip the commitment power around: the mechanism first commits to the votes of the synthetic agents, then the real agents best-respond.
This variant may be practical in some settings where the mechanism does have approximate knowledge of $\mathbf{V}$. For example, the decisionmaking process may include discussions, straw polls, etc. However, in general, we view the assumption as being an impractical one made for the sake of analysis. We view our results for the impractical variant as giving strong evidence for the performance of the Practical variant or other similar mechanisms based on QTM.
The Impractical Variant of the Synthetic Players QTM proceeds as follows.
  1. 1.
    Assume knowledge of the value totals $\mathbf{V}$ (though not individual values).
  2. 2.
    Set $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$.
  3. 3.
    Compute a solution $\mathbf{A},\mathbf{p}$ that satisfies Equation 6.
  4. 4.
    Announce $\mathbf{A}$; calculate and announce $a_{1}^{\text{mech}},\dots,a_{m}^{\text{mech}}$ from Equation 7.
  5. 5.
    Run the QTM, but increase the total votes of each alternative $k$ by $a_{k}^{\text{mech}}$ before taking the softmax. That is, collect votes $\mathbf{a}=\mathbf{a}^{1},\dots,\mathbf{a}^{n}$ from the participants, assign payments based on $\mathbf{a}$ exactly as in the QTM, and select an alternative according to $\mathbf{p}^{\prime}$ with
    $p_{k}^{\prime}=\frac{e^{a_{k}^{\text{mech}}+\sum_{i}a_{k}^{i}}}{\sum_{\ell}e^{a_{\ell}^{\text{mech}}+\sum_{i}a_{\ell}^{i}}}.$
3.2 Analysis
Focusing on the two-alternative case, we show that the Impractical Variant inherits the welfare guarantee of the QTM. We assume for equilibrium selection that if there are multiple pure-strategy equilibria, agents select the “focal” equilibrium with vote totals $\mathbf{A}$ announced by the mechanism. Denote $\operatorname{\mathbb{E}}_{k\sim p}W_{k}=\sum_{k=1}^{m}p_{k}W_{k}$ as the expected total welfare of the mechanism. In this case, our Price of Anarchy results for the QTM can be transferred to this setting.
Proposition 4.
[…]
Proof.
Budget balance follows immediately from budget balance of the QTM.
For Price of Anarchy, we first analyze a hypothetical execution of the original QTM with synthetic and real players. Then, we use properties of that game to analyze the Impractical Variant of the Synthetic Players QTM (“our mechanism”).
Given $\mathbf{v}$, let $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$. Create a set of synthetic players as follows: choose $\hat{n}\geq\frac{\max_{k}B_{k}}{2c}$ and for $j=1,\dots,\hat{n}$, define $\mathbf{\hat{v}}^{j}=\tfrac{1}{\hat{n}}\mathbf{B}$. Observe $c\geq\tfrac{1}{2}\max_{j,k}\hat{v}_{k}^{j}$.
[…]
Now, return to our mechanism. An agent $i$’s best-response depends only on $\mathbf{v}^{i}$ and on $\mathbf{a}^{\text{mech}}+\sum_{j\neq i}\mathbf{a}^{j}$, the vector of total votes excluding $i$’s. This follows because the terms of $i$’s utility that $i$ controls depends only on these quantities. So for all real agents, the utility and best responses are strategically equivalent in our mechanism and in the synthetic game where the synthetic players commit to total votes $\mathbf{a}^{\text{mech}}$. We have already shown that at least one such equilibrium $\mathbf{a}$ exists, with the vote totals $\mathbf{A}$ computed and announced by the mechanism. By the focal equilibrium assumption, the real agents play that equilibrium $\mathbf{a}$, in which case the output $\mathbf{p}$ is identical to that of the synthetic game on equilibrium $(\mathbf{a},\mathbf{a}^{\text{mech}})$. Furthermore, in every strategy profile $\mathbf{a}$ of our mechanism, social welfare $\sum_{k}p_{k}W_{k}$ is equal to the social welfare of the synthetic game in strategy profile $(\mathbf{a},\mathbf{a}^{\text{mech}})$. By Theorem 1, in any pure-strategy equilibrium of the synthetic game, including those where the synthetic players play $\mathbf{a}^{\text{mech}}$, welfare is at least $W_{1}\cdot\max\{\tfrac{1}{2},1-(2/T)^{2/5}\}$. ∎
4 Incorporating Predictions
We now consider public projects with predictions. Specifically, we propose that an information-aggregation mechanism (just called an aggregation mechanism) be operated first in order to advise the decisionmaking mechanism, which operates second. We refer to participants in the first stage as aggregators and participants in the second as decisionmakers.
Incentives for manipulation.
If the aggregators and decisionmakers are disjoint, then we can simply use any aggregation mechanism to obtain estimates of the external welfare impact $B_{k}$ of each project $k$, then use the Synthetic Players QTM previously proposed. The challenge comes when agents may participate in both mechanisms. There are two potential problems: a decisionmaker may manipulate the predictions in order to boost some alternatives’ chances of being chosen, and an aggregator may manipulate the decisionmaking mechanism in order to increase the rewards for their predictions.
4.1 General Mechanism
Formally, we model an aggregation mechanism as a game that produces predictions of the external welfare impacts of each alternative. We let $\hat{B}_{k}$ denote the produced estimate of the external welfare impact of alternative $k$, should it be chosen. The aggregation mechanisms we consider, prediction markets and wagering mechanisms, reward participants based on an observation of the actual outcome once an alternative is chosen. The combined mechanism we propose, the Synthetic players QUAdratic transfer mechanism with Predictions (SQUAP), is as follows:
  1. 1.
    Agents participate in the aggregation mechanism, which produces $\mathbf{\hat{B}}=(\hat{B}_{1},\dots,\hat{B}_{m})$.
  2. 2.
    We run the Synthetic Players QTM defined in Section 3.1 using external welfare impacts $\mathbf{\hat{B}}$. As discussed below, instead of redistributing payments, we may use them to fund the aggregation mechanism.
  3. 3.
    The aggregation mechanism observes the choice of alternative and subsequent events and rewards its participants.
Again, we do not know how to analyze SQUAP, and we turn to Impractical SQUAP, which is the same except:
  • Instead of using the Synthetic Players QTM in the decisionmaking stage, we use the Impractical Variant of the Synthetic Players QTM, which assumes knowledge of the vote totals.
  • We do not redistribute payments from the QTM.
Redistribution has the potential at least in theory to cause potential incentive problems (see Lemma 3). For simplicity, we generally assume that the mechanism keeps or “burns” the payments, or uses them to subsidize the aggregation mechanism. In practice, designers may choose to go ahead with the redistribution, as the incentive concern is likely low.
Incentive alignment.
We emphasize that any number of agents may participate in both mechanisms or in either one alone. Perhaps surprisingly, we only require two natural properties of an aggregation mechanism in order to prove welfare guarantees in equilibrium of the combined mechanism. These properties address the two possible incentive problems discussed above. First is the incentive of aggregators to manipulate the decision in order to gain higher prediction rewards.
Definition 1.
An aggregation mechanism is alternative-independent if a participant’s net expected payment does not depend on which alternative $k$ is selected by the decisionmaking mechanism.
We will see that an importance-weighting technique due to Chen et al. [2011] allows one to achieve alternative-independence, so that prediction rewards are independent of the decisionmaking mechanism (at least in expectation).
Next is the incentive of decisionmakers to manipulate the predictions $\mathbf{\hat{B}}$ in order to boost the chances of their preferred alternative(s). In general, agents can gain from manipulating the estimated external welfare even at a cost. However, we will be able to show that the decisionmaking stage’s welfare guarantees are robust to relatively small manipulations. Therefore, we only need that large manipulations are prohibitively costly. We formalize this costliness as follows. Say that a player in a game is $x$-best-responding if they can improve their net expected payoff by at most $x$ by switching to another strategy.
Definition 2.
In an aggregation mechanism, let $\mathbf{B}$ be the output of the mechanism if all agents are truthful and let $\mathbf{\hat{B}}$ be the output under strategic behavior. Given some $x>0$ and $\alpha>0$, we say the aggregation mechanism has a deviation bound $\alpha x$ if, for all strategy profiles where all participants are $x$-best-responding,
$\max_{k}|\hat{B}_{k}-B_{k}|\leq\alpha x.$
To quantify the welfare of the combined mechanism, we assume that $\mathbf{B}$, the output of the aggregation mechanism if all agents are truthful, is the true external welfare impact. Therefore, the total welfare of alternative $k$ is $W_{k}=B_{k}+V_{k}$, and we again assume $W_{1}\geq\cdots\geq W_{m}$. Because this is now a dynamic game, we also consider a small refinement of equilibrium: We say the combined mechanism is in sub-mechanism perfect pure equilibrium if it is in Nash equilibrium and, in addition, the strategies played at each subgame consisting of the Synthetic Players QTM are in pure-strategy Nash equilibrium. We also continue to assume agents play the focal pure-strategy equilibrium in the Impractical Synthetic Players QTM. In an abuse of notation, we let $\text{pPoA}(\mathbf{v},\mathbf{B})$ refer to the ratio of the welfare in the worst such sub-mechanism perfect pure equilibrium to $W_{1}$.
Theorem 2.
Define the spread of the game to be $T\coloneqq\frac{W_{1}}{max_{i,k}v_{k}^{i}}$. Suppose we run Impractical SQUAP with an aggregation mechanism that is alternative-independent and has a deviation bound $\alpha\max_{i,k}v_{k}^{i}$. Then in the two-alternative setting, there exists a choice of $c$ such that
[…]
The proof has two main steps. First, we use our strategic properties of alternative-independence and the deviation bound to show that, in equilibrium, the estimates $\mathbf{\hat{B}}$ are reasonably accurate. (Actually, their accuracy can even degrade somewhat as the spread of the game increases and pPoA will still tend to one.) Then, we modify our QTM Price of Anarchy analysis to obtain the welfare guarantee.
Lemma 3.
In any sub-mechanism perfect pure equilibrium of SQUAP, if the aggregation mechanism is alternative-independent and has a deviation bound $\alpha x$, then its output satisfies for all $k$ the bound $|\hat{B}_{k}-B_{k}|\leq\alpha\max_{i,k}v_{k}^{i}$.
Proof.
Fix any events of the aggregation mechanism and consider the subgame of the QTM. Let an agent’s net utility in the combined mechanism be $f(k)=g(k)+h(k)$ if alternative $k$ is selected, where $g(k)$ is the net payoff from the aggregation mechanism and $h(k)$ is the net utility in the QTM stage, i.e. $v_{k}^{i}-c\sum_{\ell}(a_{\ell}^{i})^{2}$. By alternative-independence, $\operatorname{\mathbb{E}}[g(k)]=C$ for some constant $C$ independent of $k$. By design of the QTM without redistribution, $\max_{k}h(k)\leq\max_{i,k}v_{k}^{i}$. It immediately follows that in any QTM subgame equilibrium, all agents are ($\max_{i,k}v_{k}^{i}$)-best-responding.
Now we aim to show that that in any combined equilibrium, in the aggregation stage all agents are ($\max_{i,k}v_{k}^{i}$)-best-responding. We proceed by contradiction: assume not. Then there exists an agent in this equilibrium who is $x$-best-responding in the aggregation mechanism for $x>\max_{i,k}v_{k}^{i}$. By definition, the agent can deviate to a strategy in the aggregation mechanism which leads to a utility increase strictly larger than $\max_{i,k}v_{k}^{i}$. Moreover, if the agent deviates to casting zero votes in the QTM, she loses at most $\max_{i,k}v_{k}^{i}$ in utility, since $f(k)-f(k^{\prime})\leq\max_{i,k}v_{k}^{i}$ for any pair of alternatives $k,k^{\prime}$. It follows that this deviation strategy leads to a strictly positive change in utility; thus, the current strategy is not an equilibrium. By contradiction, then, the statement holds.
Since all agents are $(\max_{i,k}v_{k}^{i})$-best responding, the deviation bound implies the result. ∎
We now prove Theorem 2, the Price of Anarchy bound. The general strategy mirrors our QTM analysis, where we first show that $p_{1}$ approaches $1$ for a large gap $W_{1}-W_{2}$.
Lemma 4.
If the aggregation mechanism is alternative-independent and has deviation bound $\alpha x$, then in the two alternative setting, if $W_{1}-W_{2}\geq 2\alpha\max_{i,k}v_{k}^{i}$, then in any submechanism-perfect pure equilibrium,
$p_{1}\geq 1-\left(\frac{8c}{W_{1}-W_{2}-2\alpha\max_{i,k}v_{k}^{i}}\right)^{2/3}.$
Proof.
Define $\hat{W}_{k}=V_{k}+\hat{B}_{k}$, and recall $W_{k}=V_{k}+B_{k}$. By Lemma 3, we have
$\hat{W}_{1}-\hat{W}_{2}$ $=W_{1}-W_{2}+\hat{B}_{1}-B_{1}+\hat{B}_{2}-B_{2}$
$\geq W_{1}-W_{2}-2\alpha\max_{i,k}v_{k}^{i}$
$\geq 0.$
In the QTM stage, the synthetic players’ aggregate values are $\mathbf{\hat{B}}$. Therefore, by Proposition 4, it satisfies
$p_{1}$ $\geq 1-\left(\frac{8c}{\hat{W}_{1}-\hat{W}_{2}}\right)^{2/3}$
$\geq 1-\left(\frac{8c}{W_{1}-W_{2}-2\alpha\max_{i,k}v_{k}^{i}}\right)^{2/3}.$
Proof of Theorem 2.
As with the QTM analysis, we divide into cases. Let $c=\tfrac{1}{2}\max_{i,k}v_{k}^{i}$. Let $Y$ be the magic number $(8c)^{2/5}W_{1}^{3/5}$ that balances the following two cases.
Case $W_{1}-W_{2}\leq 2\alpha\max_{i,k}v_{k}^{i}+Y$: In this case, the Price of Anarchy is at least
$\frac{W_{2}}{W_{1}}$ $=\frac{W_{1}-(W_{1}-W_{2})}{W_{1}}$
$=1-\frac{2\alpha\max_{i,k}v_{k}^{i}+Y}{W_{1}}$
$=1-\frac{2\alpha}{T}-\left(\frac{4}{T}\right)^{2/5}.$
Case $W_{1}-W_{2}\geq 2\alpha\max_{i,k}v_{k}^{i}+Y$: In this case, by Lemma 4, the Price of Anarchy is at least
$1-\left(\frac{8c}{W_{1}-W_{2}-2\alpha\max_{i,k}}\right)^{2/3}$
$=1-\left(\frac{8c}{Y}\right)^{2/3}$
$=1-\left(\frac{4}{T}\right)^{2/5}.$
In both cases, $\text{pPoA}\geq 1-\frac{2\alpha}{T}-\left(\frac{4}{T}\right)^{2/5}$. ∎
4.2 Prediction Markets
First, we consider using prediction markets to aggregate information on external welfare in the first stage. Prediction markets are financial markets designed specifically for aggregating the predictions of agents by rewarding accuracy once an outcome is observed. We briefly define scoring-rule based prediction and decision markets below, referring the reader to references in Section 1.2 for more background. For simplicity, we will suppose that agents predict $\mathbf{B}$ directly and the mechanism is later able to exactly observe $B_{k}^{*}$, the true external welfare of the alternative $k$ that was selected. One could consider markets that are more complex and predict other relevant variables, either because observing $B_{k}^{*}$ directly is difficult, or to facilitate information aggregation. In this case, the mechanism could compute an estimate of $\mathbf{B}$ from the prediction markets’ forecasts and use the estimate in SQUAP.
The market scoring rule.
We define a scoring-rule prediction market for a real-valued random variable, following e.g. Abernethy et al. [2013]. The market defines an initial estimate $\hat{b}^{0}$. Agents arrive in sequence and provide estimates $\hat{b}^{1},\dots,\hat{b}^{N}$. Later, when the outcome $b^{*}$ of the random variable is observed, the provider of each estimate $\hat{b}^{t}$ receives a net payoff $s(\hat{b}^{t},b^{*})-s(\hat{b}^{t-1},b^{*})$, where $s$ is some proper scoring rule for the mean [Savage, 1971, Gneiting and Raftery, 2007]: the expected value of $s(\hat{b},b^{*})$, over draws $b^{*}$ from some distribution $D^{*}$, is maximized by choosing $\hat{b}=\operatorname{\mathbb{E}}[b^{*}]$. In particular, here we we will focus on the quadratic score, with a parameter $\beta>0$:
$s(\hat{b},b^{*})=-\frac{1}{\beta}(\hat{b}-b^{*})^{2}.$ (9)
Observe that each participant’s net payoff can be positive or negative, and the total payment of the mechanism telescopes to $s(\hat{b}^{N},b^{*})-s(\hat{b}^{0},b^{*})$. Following the prediction market literature, we refer to $\beta$ as the liquidity parameter. $\beta$ determines how sensitive the payoffs are to small changes in a prediction. If $\beta$ is small, then the total payoffs available are very large, and a significant payment may be required for changing the prediction even a small amount.
Decision markets.
In a decision market (e.g. [Chen et al., 2011]), we simultaneously run $m$ prediction markets, one for each of $m$ alternatives. Here, we model this as each participant $t=1,\dots$ providing an estimate $\mathbf{\hat{b}}^{t}\in\operatorname{\mathbb{R}}^{m}$. Then, we make a decision $k$ based on the outcome of the markets, and later observe $b_{k}^{*}$. We cancel all trades in the other $m-1$ markets, i.e. we only assign payoffs based on predictions $\hat{b}_{k}^{t}$ made in the $k$th prediction market.
In general, decision markets can have incentives to misreport and manipulate the decision. If the decision $k$ is made from a full-support probability distribution $\mathbf{p}$, as in the QTM, then we can address this issue with the importance-weighted or inverse-weighted scoring rule proposed by Chen et al. [2011], specialized to the quadratic score:
$S(\mathbf{\hat{b}}^{t},\mathbf{p},k,b_{k}^{*})=\frac{1}{p_{k}}s(\hat{b}^{t}_{k},b_{k}^{*}).$
This rule can be impractical when probabilities are small, and practitioners may prefer to replace with it something bounded, e.g. if ex post budget balance is required. However, the theoretical benefit is that, even when $\mathbf{p}$ is chosen by some rule that depends arbitrarily on the predictions, there is no incentive to manipulate predictions to increase expected score. Specifically, if the outcome $b_{k}^{*}$ is drawn from a distribution $D_{k}^{*}$, then the expected score over both $k\sim\mathbf{p}$ and $b_{k}^{*}\sim D_{k}^{*}$ is
$\sum_{k}p_{k}\operatorname{\mathbb{E}}_{D_{k}^{*}}S(\mathbf{\hat{b}}^{t},\mathbf{p},k,b_{k}^{*})$ $=\frac{-1}{\beta}\sum_{k}\operatorname{\mathbb{E}}_{D_{k}^{*}}(\hat{b}^{t}_{k}-b_{k}^{*})^{2}$
$=\frac{-1}{\beta}\sum_{k}\left((\hat{b}^{t}_{k}-B_{k})^{2}+\text{Var}(D_{k}^{*})\right),$
where $B_{k}=\operatorname{\mathbb{E}}_{D_{k}^{*}}[b_{k}^{*}]$. The variance term is independent of any agents’ actions, and moreover cancels out in an agent’s payoff when we consider the market scoring rule (difference of two scores). Therefore, under the importance-weighted scoring rule, it is without loss of generality to define the expected score of a prediction $\mathbf{\hat{b}}$, given the vector of means $\mathbf{B}=B_{1},\dots,B_{m}$, to be
$S(\mathbf{\hat{b}},\mathbf{B})$ $=\frac{-1}{\beta}\sum_{k}(\hat{b}_{k}-B_{k})^{2}.$ (10)
Observation 1.
The decision market with the importance-weighted quadratic score is alternative-independent.
Efficient markets assumption.
Strategic behavior in prediction markets is complex and not fully understood [Chen et al., 2010, Ostrovsky, 2012, Gao et al., 2013, Chen and Waggoner, 2016]. To abstract from the details of strategic behavior, we assume that in the prediction market stage of the combined mechanism, information is fully aggregated.
Definition 3.
We say the efficient markets assumption holds on the combined mechanism if, in any equilibrium, the last participant in the decision market knows and believes $\mathbf{B}$, the true expected external welfare impacts of the alternatives conditioned on all available information.
It is known that in strategic equilibrium of prediction markets, information is always aggregated under a condition called separable securities [Ostrovsky, 2012]. In other words, when information aggregation fails in prediction markets, it is not for strategic reasons, but only because of the expressiveness of the prediction language. Therefore, we believe that efficient markets is a mild assumption on strategic behavior of the agents in the combined mechanism. However, it would not hold when, for example, a participant is the sole holder of information that alternative $k$ has very low welfare, and that participant has a high personal value for choosing $k$.
4.2.1 Price of Anarchy
We can now analyze Impractical SQUAP with a decision market aggregation mechanism.
Lemma 5.
Under the efficient market assumption, for any $x>0$, the importance-weighted decision market with the quadratic scoring rule and liquidity parameter $\beta=\epsilon x$ satisfies a deviation bound of $\epsilon^{1/2}x$.
Proof.
Observe that under the efficient market assumption, the decision market is equivalent to the following: run the entire market, where the last participant predicts $\mathbf{B}$. Then, give the last participant one more opportunity to participate; denote their final predictions by $\mathbf{\hat{B}}$. This is equivalent because of the telescoping sum of the market scoring rule, so that the two prediction opportunities have the same net payoff as one opportunity with predicting $\mathbf{\hat{B}}$.
For any participant except the final one, under the efficient market assumption, their strategy does not affect the final predictions $\hat{B}$. One way to see this is by the above equivalence, since the market will end at $\mathbf{B}$ regardless, followed by the last participant’s possible manipulation.
So we only need to analyze the incentives of the final participant $i$. By the decomposition above, we only need to consider the incentives for the final prediction opportunity (since the previous one is fixed to $\mathbf{B}$). The expected utility for $\mathbf{\hat{B}}$ is
$S(\mathbf{\hat{B}};\mathbf{B})-S(\mathbf{B};\mathbf{B})$ $=S(\mathbf{\hat{B}};\mathbf{B})$
$=-\frac{1}{\beta}\sum_{k}(\hat{B}_{k}-B_{k})^{2}.$
By the definition of $x$-best-responding in the market, $x>\frac{1}{\beta}\sum_{k}(\hat{B}_{k}-B_{k})^{2}$. So for all $k$, $(\hat{B}_{k}-B_{k})^{2}\leq\beta x=\epsilon x^{2}$, or $|\hat{B}_{k}-B_{k}|\leq\sqrt{\epsilon}x$. ∎
From Theorem 2, we immediately get the following.
Corollary 3.
Let $T=\frac{W_{1}}{\max_{i,k}v_{k}^{i}}$ be the spread. Under the efficient markets assumption, Impractical SQUAP with market parameter $\beta=\epsilon\max_{i,k}v_{k}^{i}$ and QTM parameter $c=\tfrac{1}{2}\max_{i,k}v_{k}^{i}$ satisfies
[…]
We observe that, even with constant liquidity $\epsilon=\Theta(1)$, the Price of Anarchy tends to one with the spread. Typically in a prediction market, we have $\epsilon\to 0$ with the size of the market [Abernethy et al., 2014]. We would expect $\epsilon\to 0$ in large settings in practice, but it is not required for good Price of Anarchy. This occurs because the QTM is robust to larger and larger approximations in $\mathbf{W}$ as the spread of the market grows. While one might prefer a larger decision market in practice in hopes that it better aggregates information, this robustness may at least be reassuring. We will also utilize it next to obtain budget balance in many cases.
4.2.2 Budget balance
Prediction markets are typically assumed to be subsidized. Although the subsidy could come from charging transaction fees in the market, a natural question is whether the transfers of the QTM portion of the mechanism can be used to fund the aggregation mechanism. In general, this is not always achievable, because as shown by Eguia et al. [2019, 2023], the revenue of the QTM stage can shrink to zero even as the spread of the market grows. However, using Proposition 2, we can characterize settings with growing revenue and that therefore allow for a budget balanced mechanism that self-funds its decision market.
Corollary 4.
[…]
Proof.
We prove that the mechanism can choose liquidity parameters $\beta$ so that the expected revenue is larger than the amount spent on the decision market stage in expectation. In practice, one would like to scale the prediction-market payments by the largest possible amount so that budget balance is achieved. However, we cannot rule out manipulation for reasons similar to the redistribution, so for analysis we assume that the mechanism designer commits to $\beta$ up front. By the telescoping nature of the proper scoring rule, the mechanism’s total spend on the decision market stage in expectation is $S(\mathbf{\hat{B}};\mathbf{B})-S(\mathbf{B^{0}},\mathbf{B})\leq S(\mathbf{B};\mathbf{B})-S(\mathbf{B^{0}};\mathbf{B})=\frac{1}{\beta}\sum_{k}(B_{k}^{0}-B_{k})^{2}$. With at-least-constant disagreement, revenue is at least constant, so $\beta=\Omega(1)$ is possible with ex ante budget-balance. By Corollary 3, in this case and with $T\to\infty$, the Price of Anarchy converges to $1$. ∎
4.3 Wagering Mechanisms
Wagering mechanisms are another common tool for eliciting and aggregating predictions [Lambert et al., 2008, 2015b]. These mechanisms can be preferable to prediction markets because they are one-shot, and thus simpler to implement; moreover, they are budget balanced. However, whereas prediction markets automatically aggregate information, in a wagering mechanism the designer must combine the resulting predictions from a wagering mechanism herself (often by averaging).
We consider a natural extension that does not appear to have been studied: decision wagering mechanisms, analogous to decision markets. Instead of agents wagering on a single future event, we will have them wager on the future event conditional on alternative $k$ being selected, for each $k=1,\dots,m$. Once $k$ is selected and the event observed, we cancel all other wagers and assign payments based on the predictions for $k$.
Specifically, we adapt the Brier betting mechanism [Lambert et al., 2015a] to our setting, defining the importance-weighted quadratic decision wagering mechanism. For each alternative $k$, forecaster $i$ submits an estimate $\hat{b}^{i}_{k}$ for the expected external welfare conditioned on that alternative. Let $\mathbf{\hat{b}}$ consist of all the players’ predictions for all the alternatives. Once an alternative $k\sim\mathbf{p}$ is chosen and the random variable $b^{*}_{k}$ is observed, each player $i$ receives the payoff
$\pi_{i}(\mathbf{\hat{b}},\mathbf{p},k,b_{k}^{*})=\frac{1}{p_{k}}\left[s(\hat{b}^{i}_{k},b^{*}_{k})-\frac{1}{N}\sum_{j=1}^{N}s(\hat{b}^{j}_{k},b^{*}_{k})\right],$
where $s$ is the quadratic score defined in Equation 9. In other words, forecaster $i$’s payoff consists of her score minus the average score of all participants. Since $\sum_{i}\pi_{i}(\mathbf{\hat{b}},\mathbf{p},k,b_{k}^{*})=0$, we obtain:
Observation 2.
The importance-weighted quadratic decision wagering mechanism achieves budget-balance.
Wagers and agent loss.
Above, we have considered the special case of a wagering mechanism where all wagers are equal to one. In general, the wagering mechanism weights each participant’s score $s(\hat{b}^{i}_{k},b^{*}_{k})$ by $\frac{y_{i}^{k}}{Y_{k}}$, where $y^{i}_{k}$ is some nonnegative wager on alternative $k$ and $Y_{k}=\sum_{i}y^{i}_{k}$. The wagering mechanism is generally applied in cases where scores $s()$ are bounded in $[0,1]$, so that an agent can be guaranteed to win or lose no more than her wager. In our case, such guarantees can be achieved by adjusting $\beta$, the liquidity parameter of $s$, if an upper bound on $B_{k}^{*}$ is known. Wagering mechanisms can also be subsidized, e.g. by additionally giving each agent a fraction of their score, or by making all payouts nonnegative with a shift. We could use a similar analysis to the prediction market setting to use the QTM revenue to subsidize the wagering mechanism.
Strategic behavior and aggregation.
We adopt the standard wagering mechanism model of immutable beliefs, where each agent $i$ has fixed beliefs $\mathbf{b}^{i}$ about the expectation of $\mathbf{b}^{*}$, regardless of others’ beliefs. (In a Bayesian model, we essentially have a “no-trade” situation in which the zero-sum payoffs discourage participation.) The wagering mechanism in isolation is strategyproof: agents maximize expected net payoff by reporting their true believed estimates. This follows immediately from the properness of the scoring rule $s$ and the fact that in $\pi_{i}$, the agent can only affect their own score. For analysis, we suppose that the wagering mechanism outputs $\hat{B}_{1},\dots,\hat{B}_{m}$ where $\hat{B}_{k}=\frac{1}{N}\sum_{i=1}^{N}\hat{b}_{k}^{i}$. We assume that, if all agents are truthful, then the output $\mathbf{B}$ is the true expectation of the external welfare impacts. Our analysis could be adapted to other aggregation methods than taking the average.
4.3.1 pPoA guarantee
It only remains to show the deviation bound. Because the wagering mechanism relies on the same quadratic scoring rule as the prediction market, the analysis (Appendix B) uses the same tools, but is simpler.
Lemma 6.
The importance-weighted quadratic decision wagering mechanism with parameter $\beta=\epsilon x$ has a deviation bound $\epsilon^{1/2}x$.
Corollary 5.
Let $T=\frac{W_{1}}{\max_{i,k}v_{k}^{i}}$ be the spread. Impractical SQUAP with the importance-weighted quadratic decision mechanism with wagering parameter $\beta=\epsilon\max_{i,k}v_{k}^{i}$ and QTM parameter $c=\tfrac{1}{2}\max_{i,k}v_{k}^{i}$ satisfies
[…]
5 Discussion
Summary.
This paper’s first contribution was a Price of Anarchy analysis of the Quadratic Transfers Mechanism, primarily in the two-alternative setting. Prior work had welfare analysis in asymptotic settings with randomly drawn agents. But results for such a setting, reminiscent of a “strategic jury theorem”, left open a question of robust guarantees with arbitrary sets of agents, as we gave in Section 2.
The second contribution was to extend the QTM to a case with external welfare impacts of the decisions. Many public projects settings involve impacts on non-decisionmakers, such as the climate impacts of a particular policy. Section 3 proposed a practical mechanism for this problem, the Synthetic Players QTM. We were able to show strong Price of Anarchy guarantees for a less-practical variant in which the decisionmaker needs to estimate the total values of each alternative. It approximately maximizes the welfare of decisionmakers plus externalities of the decision.
The third and main contribution was to propose decisionmaking mechanisms that combine preference aggregation with information aggregation. We model the information aggregation stage as forecasting of the external welfare impacts of the decision. By using either prediction markets or wagering mechanisms, we again showed that a less-practical variant of the Synthetic Players QTM with Predictions (SQUAP) has strong Price of Anarchy guarantees. The importance-weighting technique gave forecasters the same expected payoff for any decision of the mechanism. The proper scoring rules enforced penalties for inaccurate predictions, which can only change the final decision with a small probability while causing a prohibitive penalty.
Future work.
There are a number of technical open problems. First is improving the QTM Price of Anarchy guarantee for more than two alternatives, or giving counterexamples for parameter regimes where it is impossible. Second is proving a Price of Anarchy guarantee for the Synthetic Players QTM (i.e. the practical variant), or finding a counterexample. Even in the latter case, we expect a “Price of Stability” result to be possible. Third is to extend SQUAP to other information-aggregation mechanisms, perhaps with a more practical method than importance weighting, and prove guarantees for the practical variant.
Conceptually, there are many other possible approaches to decisionmaking with preferences and predictions. In particular, our approach used monetary mechanisms and the quantitative criterion of social welfare maximization, which required a strong quasilinear assumption on participants. Even within that sphere, we have only proposed one possible approach. But outside of it, many other directions are possible.
Acknowledgements
Supported by the Ethereum Foundation grant FY22-0716.
Appendix A Quadratic Transfer Mechanism
This section gives proofs from Section 2 along with some additional results and discussion.
A.1 Equilibrium Characterization
This section contains proofs and discussion from Section 2.2
To begin, recall that a strategy is strictly dominated if there exists another strategy that always obtains strictly higher utility.
Lemma 7.
For all agents $i$, if the pure strategy $\mathbf{a}^{i}$ has $|a_{k}^{i}|>\sqrt{\frac{max_{l}v_{l}^{i}}{c}}$ for any $k$, then $\mathbf{a}^{i}$ is a strictly dominated strategy.
Proof.
Fix all other agents’ strategies as $\mathbf{a}^{-i}$. Denote the expected utility $u^{i}$ of pure strategy $\mathbf{a}^{i}$ given fixed strategies $\mathbf{a}^{-i}$ as $u^{i}(\mathbf{a}^{i};\mathbf{a}^{-i})$, and the softmax probability $p_{k}$ over strategy $a^{i}$ and fixed strategies $\mathbf{a}^{-i}$ as $p_{k}(\mathbf{a}^{i};\mathbf{a}^{-i})$. Consider the strategy $\hat{a}_{k}^{i}=0\;\forall k\in\mathbb{Y}$. Then
$u^{i}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})$ $=\sum_{k=1}^{m}p_{k}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})v_{k}^{i}-c\sum_{k=1}^{m}0^{2}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$ (13)
$\geq\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}.$ (16)
Now, take a different pure strategy $\mathbf{\tilde{a}}^{i}$ for voter $i$ where $|\tilde{a}_{k}^{i}|>\sqrt{\frac{max_{l}v_{l}^{i}}{c}}$ for some $k\in\mathbb{Y}$. Since $(\tilde{a}_{k}^{i})^{2}$ is nonnegative, we must have $\sum_{k=1}^{m}(\tilde{a}_{k}^{i})^{2}>\frac{1}{c}(\max_{l}v_{l}^{i})$. Thus
$u^{i}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})$ $=\sum_{k=1}^{m}p_{k}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})v_{k}^{i}-c\sum_{k=1}^{m}(\mathbf{\tilde{a}}_{k}^{i})^{2}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$ (19)
$<\max_{l}v_{l}^{i}-\max_{l}v_{l}^{i}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$ (22)
$=\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}.$ (25)
Combining Inequalities 16 and 25, for any strategy profile $\mathbf{a}^{-i}$ of voters other than $i$, $u^{i}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})<\frac{1}{n-1}\sum_{j\neq i}\sum_{k=1}^{m}(a_{k}^{j})^{2}<u^{i}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})$. By definition, then, it is a strictly dominated strategy for any agent $i$ to play $\mathbf{\tilde{a}}^{i}$. ∎
Lemma 8.
Fix a set of agents with types $\mathbf{v}$. Let QTM’ be the QTM where each agent is restricted to the compact strategy space $\Bigl{[}-\sqrt{\frac{\max_{i,k}v_{k}^{i}}{c}},\sqrt{\frac{\max_{i,k}v_{k}^{i}}{c}}\Bigr{]}^{m}$. A pure-strategy Nash equilibrium exists for QTM’ if and only if that same pure-strategy equilibrium exists for QTM.
Proof.
Let $\mathcal{A}:=\Bigl{[}-\sqrt{\frac{\max_{i,k}v_{k}^{i}}{c}},\sqrt{\frac{\max_{i,k}v_{k}^{i}}{c}}\Bigr{]}^{m}$. Suppose $\mathbf{a}_{*}=(\mathbf{a}^{1}_{*},\mathbf{a}^{2}_{*},...\mathbf{a}^{n}_{*})$ is a pure-strategy equilibrium for QTM’. Then for each $i$, $\mathbf{a}_{*}^{i}$ is a best response among $\mathcal{A}$ to $\mathbf{a}_{*}^{-i}$. By Lemma 7, no strategy $\mathbf{a}^{i}\not\in\mathcal{A}$ can be a best response to $\mathbf{a}_{*}^{-i}$, because $\mathbf{a}^{i}$ is strictly dominated by some strategy in $\mathcal{A}$. So $\mathbf{a}_{*}$ remains an equilibrium for QTM.
Conversely, let $\mathbf{a}_{*}$ be an equilibrium for QTM. We must have $\mathbf{a}_{*}^{i}\in\mathcal{A}$ for all $i$, as strictly dominated strategies are never best responses. So $\mathbf{a}_{*}$ is an equilibrium for QTM’, as the strategies all lie in $\mathcal{A}$ and are all best responses among all strategies in $\mathcal{A}$. ∎
Recall that Lemma 1 stated that if $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then agent utilities are strictly concave on $\operatorname{\mathbb{R}}^{m}$.
Proof of Lemma 1.
Note that $\hat{u}^{i}=u^{i}/c$ has the same concavity as $u^{i}$. Consider $H^{i}$, the Hessian matrix of $\hat{u}^{i}$. We begin by computing expressions for diagonal entries $H_{kk}^{i}$ and non-diagonal entries $H_{kl}$, where $k\neq l$.
$\frac{\partial\hat{u}^{i}}{\partial a_{k}^{i}}$ $=\sum_{h=1}^{m}\frac{\partial}{\partial a_{k}^{i}}p_{h}\frac{v_{h}^{i}}{c}-2a_{k}^{i},\text{so}$
$H_{kk}^{i}=\frac{\partial^{2}\hat{u}^{i}}{\partial a_{k}^{i}\partial a_{k}^{i}}$ $=\sum_{h=1}^{m}\frac{\partial^{2}}{(\partial a_{k}^{i})^{2}}p_{h}\frac{v_{h}^{i}}{c}-2$ (26)
$H_{kl}^{i}=\frac{\partial^{2}\hat{u}^{i}}{\partial a_{k}^{i}\partial a_{l}^{i}}$ $=\sum_{h=1}^{m}\frac{\partial^{2}}{\partial a_{k}^{i}\partial a_{l}^{i}}p_{h}\frac{v_{h}^{i}}{c}.$ (27)
The first derivatives for each probability $p_{k}$ are listed below:
$\frac{\partial p_{k}}{\partial a_{k}^{i}}$ $=\frac{e^{A_{k}}\sum_{h=1}^{m}e^{A_{h}}-e^{2A_{k}}}{\Bigl{(}\sum_{h=1}^{m}e^{A_{h}}\Bigr{)}^{2}}=p_{k}-p_{k}^{2}$
$\frac{\partial p_{l}}{\partial a_{k}^{i}}$ $=\frac{-e^{A_{l}}e^{A_{k}}}{\Bigl{(}\sum_{h=1}^{m}e^{A_{h}}\Bigr{)}^{2}}=-p_{l}p_{k}.$
We extend to the second derivatives as well; given $k\neq l\neq h$,
$\frac{\partial^{2}p_{k}}{\partial a_{k}^{i}\partial a_{k}^{i}}$ $=p_{k}(1-3p_{k}+2p_{k}^{2})=p_{k}(2p_{k}-1)(p_{k}-1)$
$\frac{\partial^{2}p_{k}}{\partial a_{k}^{i}\partial a_{l}^{i}}$ $=p_{k}p_{l}(2p_{k}-1)$
$\frac{\partial^{2}p_{k}}{\partial a_{l}^{i}\partial a_{h}^{i}}$ $=2p_{l}p_{h}p_{k}$
$\frac{\partial^{2}p_{k}}{\partial a_{l}^{i}\partial a_{l}^{i}}$ $=p_{k}p_{l}(2p_{l}-1).$
Denote $\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i}=\sum_{k=1}^{m}p_{k}v_{k}^{i}$, the expected value of the mechanism for agent $i$. Plugging the above values into Equations (26) and (27), we have
$H_{kk}^{i}$ $=\sum_{h=1}^{m}\frac{\partial^{2}}{(\partial a_{k}^{i})^{2}}p_{h}\frac{v_{h}^{i}}{c}-2$
$=\frac{p_{k}}{c}\Bigl{[}(2p_{k}-1)(p_{k}-1)v_{k}^{i}+(2p_{k}-1)\sum_{h\neq k}p_{h}v_{h}^{i}\Bigr{]}-2$
$=\frac{p_{k}}{c}(2p_{k}-1)(\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i}-v_{k}^{i})-2,$
$H_{kl}^{i}$ $=\sum_{h=1}^{m}\frac{\partial^{2}}{\partial a_{k}^{i}\partial a_{l}^{i}}p_{h}\frac{v_{h}^{i}}{c}$
$=\frac{1}{c}\Bigl{[}p_{k}p_{l}(2p_{k}-1)v_{k}^{i}+p_{k}p_{l}(2p_{l}-1)v_{l}^{i}+\sum_{h\neq k,l}2p_{l}p_{h}p_{k}v_{h}^{i}\Bigr{]}$
$=\frac{p_{k}p_{l}}{c}(2\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i}-v_{k}^{i}-v_{l}^{i}).$
We consider the matrix $B^{i}$ such that $H^{i}=-2I+B^{i}$. Then the inequality $v^{T}H^{i}v<0$ corresponds to $v^{T}B^{i}v<2$: so to prove $H^{i}$ is negative definite, we just need to bound the spectral radius of $B^{i}$ by 2.
As an interesting side note, the sum over column $k$ in $B^{i}$ is
$\biggl{(}\sum_{l\neq k}\frac{\partial^{2}\hat{u}^{i}}{\partial a_{k}^{i}\partial a_{l}^{i}}\biggr{)}+\frac{\partial^{2}\hat{u}^{i}}{\partial a_{k}^{i}\partial a_{k}^{i}}+2$ $=\biggl{(}\sum_{l\neq k}\frac{p_{k}p_{l}}{c}(2\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i}-v_{k}^{i}-v_{l}^{i})\biggr{)}+\frac{p_{k}}{c}(2p_{k}-1)(\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i}-v_{k}^{i})=0.$
Now, the spectral radius of a matrix is bounded by the maximum sum of absolute value of entries in any row (in our case, $\sum_{l=1}^{m}|B^{i}_{kl}|$). Note that $|B_{kk}^{i}|=|H_{kk}^{i}+2|\leq\tfrac{p_{k}}{c}(2p_{k}-1)\max_{k}{v^{i}_{k}}$ and $|B_{kl}|=|B_{lk}^{i}|=|H_{kl}^{i}|\leq(2p_{k}p_{l}\max_{k}{v^{i}_{k}})/c$. So
$\sum_{l=1}^{m}|B_{kl}^{i}|$ $\leq\frac{2p_{k}\max_{k}{v^{i}_{k}}}{c}\sum_{l\neq k}p_{l}+p_{k}(2p_{k}-1)\frac{\max_{k}{v^{i}_{k}}}{c}$
$=\frac{\max_{h}{v^{i}_{h}}}{c}(2p_{k}(1-p_{k})+p_{k}(2p_{k}-1))$
$=\frac{p_{k}\max_{h}{v^{i}_{h}}}{c}<\frac{\max_{h}{v^{i}_{h}}}{c}.$
Thus, for $c\geq\frac{\max_{h}{v^{i}_{h}}}{2}$, the maximum sum of the absolute value of entries in any row of $B^{i}$ is strictly less than 2; the spectral radius of $B^{i}$ is strictly less than 2; and $H^{i}$ is negative definite for this value of $c$, meaning that $u^{i}$ is strictly concave. ∎
Remark 1.
The above analysis can be reinterpreted in terms of exponential family distributions. Specifically, the mechanism generates probabilities representing which one of $m$ events $X_{1},X_{2},\ldots,X_{m}$ will occur, and thus corresponds to a multinomial distribution with parameter $\mathbf{p}$. An exponential family distribution’s $n$th cumulant can be calculated directly as the $n$th derivative of the cumulant function $A$. One can verify that the derivative of $A$ with respect to the $k$th parameter corresponds to $p_{k}$, the mean of the random variable $X_{k}$. It follows that the first partial derivatives of $p$ correspond to the variance and covariance, and the second partial derivatives correspond to the third cumulant.
A.2 Two Alternatives
This section contains proofs from Section 2.3.
In the rest of this section, we prove Theorem 1,. We first observe an equivalent representation of the first-order conditions in the two-alternative case.
Corollary 6.
In the two-alternative case, in equilibrium we have $A_{1}=\frac{p_{1}p_{2}}{2c}(V_{1}-V_{2})$.
Proof.
By Lemma 2,
$a_{1}^{i}$ $=\frac{p_{1}}{2c}(v_{1}^{i}-\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i})$
$=\frac{p_{1}}{2c}(v_{1}^{i}-p_{1}v_{1}^{i}-(1-p_{1})v_{2}^{i})$
$=\frac{p_{1}p_{2}}{2c}(v_{1}^{i}-v_{2}^{i}).$
It follows that
$A_{1}=\frac{p_{1}p_{2}}{2c}(V_{1}-V_{2}).$
Our first result is that the Price of Anarchy in the two-alternative case always exceeds $1/2$. This follows because, by Corollary 6 and $V_{1}\geq V_{2}$, there are always at least as many votes for alternative $1$ as for $2$.
Proposition 5.
[…]
Proof.
$p_{1}=\frac{e^{A_{1}}}{e^{A_{1}}+e^{A_{2}}}$ and $p_{2}=\frac{e^{A_{2}}}{e^{A_{1}}+e^{A_{2}}}$. By Lemma 2, $A_{1}=\frac{p_{1}p_{2}}{2c}(V_{1}-V_{2})\geq 0$. Also by Lemma 2, $A_{2}=-A_{1}\leq 0$. Thus $p_{1}\geq p_{2}$, so $p_{1}\geq 1/2$. If $V_{1}>V_{2}$, then $A_{1}>0$ and $p_{1}>p_{2}$. If $V_{1}=V_{2}$ then $\text{pPoA}=1$; otherwise, $\text{pPoA}(\mathbf{v})\geq p_{1}>1/2$. ∎
We now consider the case where there is an additive gap between the aggregate values of the alternatives. This is the main result behind both Theorem 1 and Corollary 2.
Lemma 9.
In any pure-strategy equilibrium of the two-alternative QTM,
$p_{1}\geq 1-\left(\frac{8c}{V_{1}-V_{2}}\right)^{2/3}.$
Proof.
Because votes sum to zero, $A_{1}=-A_{2}$, and we can write $p_{2}=e^{-A_{1}}/(e^{A_{1}}+e^{-A_{1}})$. Furthermore, By Corollary 6, in equilibrium we have the following, and use the observation that $A_{1}\geq 0$:
$A_{1}$ $=\frac{p_{1}p_{2}}{2c}(V_{1}-V_{2})$
$=\frac{V_{1}-V_{2}}{2c\left(e^{A_{1}}+e^{-A_{1}}\right)^{2}}$
$\geq\frac{V_{1}-V_{2}}{2c\left(2e^{A_{1}}\right)^{2}}$
$\implies\quad A_{1}e^{2A_{1}}$ $\geq\frac{V_{1}-V_{2}}{8c}$ (28)
$\implies\quad e^{A_{1}}$ $\geq\left(\frac{V_{1}-V_{2}}{8c}\right)^{1/3},$ (29)
where the final inequality follows because $A_{1}\leq e^{A_{1}}$, after which we cube-root both sides.
Now, if $x:=e^{A_{1}}$, then $p_{2}=\frac{1/x}{x+1/x}=\frac{1}{x^{2}+1}\leq\frac{1}{x^{2}}$. By (29), $p_{2}\leq\left(\frac{8c}{V_{1}-V_{2}}\right)^{2/3}$, proving the bound. ∎
Remark 2.
An asymptotically better convergence rate than in Theorem 1 and Corollary 2 is achievable. By (28), for all $\alpha>0$, $e^{A_{1}(2+\alpha)}=\Omega\left(\frac{V_{1}-V_{2}}{8c}\right)$. Following the same logic as Lemma 9 and Corollary 2, we obtain that for all $\epsilon>0$, $\text{pPoA}(\mathbf{v})\geq 1-O\left(\left(\frac{1}{G}\right)^{1-\epsilon}\right)$. Following the same logic as in the proof of Theorem 1, it will turn out that the factor $\left(\frac{2}{T}\right)^{2/5}$ can be improved to $O\left(\frac{2}{T}\right)^{\frac{1}{2}-\epsilon}$ for any $\epsilon>0$, at the cost of a correspondingly large constant factor.
Proof of Theorem 1.
Let $c=\tfrac{1}{2}\max_{i,k}v_{k}^{i}$. In this case, by Proposition 1, a pure-strategy equilibrium is guaranteed to exist. Let $Y=(8c)^{2/5}(2V_{1})^{3/5}$, a magically chosen quantity that balances the following cases.
Case $V_{1}-V_{2}\leq Y$: here, we observe that the Price of Anarchy is
$\frac{p_{1}V_{1}+p_{2}V_{2}}{V_{1}}$ $\geq\frac{V_{1}+V_{2}}{2V_{1}}$ using $p_{1}\geq\tfrac{1}{2}$
$\geq\frac{2V_{1}-Y}{2V_{1}}$
$=1-\frac{Y}{2V_{1}}$
$=1-\left(\frac{4c}{V_{1}}\right)^{2/5}.$
Case $V_{1}-V_{2}\geq Y$: here, we apply Lemma 9 to obtain that the Price of Anarchy exceeds
$p_{1}$ $\geq 1-\left(\frac{8c}{Y}\right)^{2/3}$
$=1-\left(\frac{4c}{V_{1}}\right)^{2/5}.$
In both cases, we obtain $1-\left(\frac{2}{T}\right)^{2/5}$. ∎
Proof of Proposition 2.
We heavily use that, in equilibrium, each individual’s votes sum to zero.
By the first-order conditions , $A_{1}=\frac{p_{1}p_{2}}{2c}(V_{1}-V_{2})$, or rearranging,
$p_{1}p_{2}$ $=\frac{2cA_{1}}{V_{1}-V_{2}}.$ (30)
Also by , $a_{1}^{i}=\frac{p_{1}p_{2}}{2c}(v_{1}^{i}-v_{2}^{i})$. So the revenue is
$\sum_{i}\left[c(a_{1}^{i})^{2}+c(a_{2}^{i})^{2}\right]$ $=2c\sum_{i}(a_{1}^{i})^{2}$
$=2c\left(\frac{p_{1}p_{2}}{2c}\right)^{2}\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}$ (31)
$=2c\left(\frac{A_{1}}{V_{1}-V_{2}}\right)^{2}\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}$
$=2c(A_{1})^{2}\frac{\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}}{(V_{1}-V_{2})^{2}}.$ (32)
It only remains to bound $(A_{1})^{2}$.
$p_{1}p_{2}$ $=\frac{e^{A_{1}}e^{-A_{1}}}{(e^{A_{1}}+e^{-A_{1}})^{2}}$
$=\frac{1}{(e^{A_{1}}+e^{-A_{1}})^{2}}$
$\geq\frac{1}{(2e^{A_{1}})^{2}}$ (33)
$=\frac{1}{4e^{2A_{1}}}.$
So
$A_{1}$ $=\frac{p_{1}p_{2}}{2c}(V_{1}-V_{2})$
$\geq\frac{V_{1}-V_{2}}{8ce^{2A_{1}}}.$
Rearranging and taking the logarithm of both sides,
$\ln\left(\frac{V_{1}-V_{2}}{8c}\right)$ $\leq 2A_{1}+\ln(A_{1})$
$\leq 3A_{1}.$
It follows that
$A_{1}^{2}$ $\geq\left(\max\left\{0~{},~{}\frac{1}{3}\ln\tfrac{V_{1}-V_{2}}{8c}\right\}\right)^{2}.$
On the other hand, we can replace Inequality 33 as follows:
$p_{1}p_{2}$ $=\frac{1}{(e^{A_{1}}+e^{-A_{1}})^{2}}$
$\leq\frac{1}{(e^{A_{1}})^{2}},$
and continuing with almost the same analysis, using $A_{1}\geq 0$ we obtain
$\ln\left(\frac{V_{1}-V_{2}}{2c}\right)$ $\geq 2A_{1}.$
We have obtained
$\max\left\{0~{},~{}\frac{1}{3}\ln\left(\frac{V_{1}-V_{2}}{8c}\right)\right\}\leq A_{1}\leq\frac{1}{2}\ln\left(\frac{V_{1}-V_{2}}{2c}\right),$
so
$\frac{2c}{9}\frac{\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}}{(V_{1}-V_{2})^{2}}\left(\max\left\{0~{},~{}\ln\tfrac{V_{1}-V_{2}}{8c}\right\}\right)^{2}\leq\text{revenue}\leq\frac{c}{2}\frac{\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}}{(V_{1}-V_{2})^{2}}\left(\ln\tfrac{V_{1}-V_{2}}{2c}\right)^{2}.$
A.3 QTM with m alternatives
Here we prove Proposition 3, that $\text{pPoA}\geq\frac{1}{m}$ for the QTM.
Proof of Proposition 3.
By Proposition 1, for this regime of $c$, a pure-strategy equilibrium exists. Assume WLOG that $V_{1}\geq\cdots\geq V_{m}$. Recall that $\mathbf{p}$ is the mechanism’s output, i.e. distribution on alternatives, and the welfare is $\sum_{\ell}p_{\ell}V_{\ell}$. Let
$\bar{V}\coloneq\sum_{\ell}p_{\ell}V_{\ell},$
the welfare of the mechanism. We seek to prove $\frac{\bar{V}}{V_{1}}\geq\frac{1}{m}$. We will lean on the first-order conditions for equilibrium (Lemma 2): reproducing Equation (5), for each alternative $k$,
$A_{k}=\frac{p_{k}}{2c}\left(V_{k}-\bar{V}\right).$ (34)
We will consider the following subsets of the alternatives:
  • $S=\{k:V_{k}\geq\bar{V}\}$.
  • $T=\{k:V_{k}<\bar{V}\}$.
  • $X=\{1\}\cup T$.
  • $Y=S\setminus\{1\}$.
Observe that $1\in S$ and $m\in T$ by Equation 34. Let $p(S)=\sum_{k\in S}p_{k}$, the probability of an outcome in $S$, and similarly for $p(T),p(X),p(Y)$. Let $w(S)=\frac{1}{p(S)}\sum_{k\in S}p_{k}V_{k}$, the expected welfare conditioned on picking an outcome in $S$, and similarly for $w(T),w(X),w(Y)$. We will deal with the case $|Y|=0$ separately.
We first claim $\frac{w(X)}{p(X)}\geq\frac{1}{m}$, or in other words, the expected welfare conditioned on being in $X$ exceeds $\frac{V_{1}}{m}$. By (34), $A_{1}>0$ and $A_{k}<0$ for all $k\in T$. Therefore, $e^{A_{1}}>e^{A_{k}}$ for all $k\in T$. Therefore,
$w(X)$ $=\frac{1}{p(X)}\left(p_{1}V_{1}+\sum_{k\in T}p_{k}V_{k}\right)$
$\geq\frac{p_{1}}{p(X)}V_{1}$
$=\frac{e^{A_{1}}}{e^{A_{1}}+\sum_{k\in T}e^{A_{k}}}V_{1}$
$\geq\frac{e^{A_{1}}}{e^{A_{1}}(|T|+1)}V_{1}$
$=\frac{V_{1}}{|T|+1}$
$\geq\frac{V_{1}}{m}.$
Now, if $|Y|=0$, then $|X|=m$ and $w(X)=\text{Welfare}$, and we are done. Otherwise, we claim $w(Y)\geq\bar{V}$, because for all $k\in Y$, since $k\in S$, we have $V_{k}\geq\bar{V}$. But since $Y$ and $Z$ partition the alternatives,
$\bar{V}$ $=p(Y)w(Y)+p(Z)w(Z)$
$\geq p(Y)\frac{V_{1}}{m}+p(Z)\bar{V}$
$\implies\bar{V}\geq\frac{V_{1}}{m}.$
Appendix B Wagering Mechanisms
Lemma (Lemma 6 restated).
The importance-weighted quadratic decision wagering mechanism with parameter $\beta=\epsilon x$ has a deviation bound $\epsilon^{1/2}x$.
Proof of Lemma 6.
Consider any participant $i$. The only terms in their expected utility that $i$ controls are $\frac{1}{p_{k}}s(\hat{b}_{k}^{i},b_{k}^{*})$ for each $k$. Exactly as in the prediction-market analysis , the expected utility of this term is given by Equation 10, $S(\mathbf{\hat{b}}^{i},\mathbf{b}^{i})$, where $\mathbf{b}^{i}$ is their belief. (The same comments about the variance term apply.) By the same arguments as in Lemma 5, with parameter $\beta=\epsilon x$, the loss from misreporting is
$S(\mathbf{\hat{b}}^{i},\mathbf{b}^{i})-S(\mathbf{b}^{i},\mathbf{b}^{i})=\frac{-1}{\epsilon x}\sum_{k}(\hat{b}^{i}_{k}-b^{i}_{k})^{2}.$
If the loss is bounded by $x$, then $|\hat{b}^{i}_{k}-b^{i}_{k}|\leq\epsilon^{1/2}x$.
Now, this holds for all participants, and the output of the wagering mechanism is $\hat{B}_{k}=\frac{1}{N}\sum_{i=1}^{N}\hat{b}^{i}_{k}$, so the mechanism has a deviation bound $\epsilon^{1/2}x$. ∎